OpenStudy (rayep12):

Write the equation of a line that is perpendicular to y=-0.3x+6 and that passes through the point (3, -8)

OpenStudy (rayep12):

@mathmate @sammixboo @TheSmartOne

OpenStudy (rayep12):

@.Sam.

OpenStudy (amruta1202):

Line is perpendicular to the given line means its slope is -1/m , if we consider m as the slope of given line

TheSmartOne (thesmartone):

perpendicular lines have the opposite reciprocal slopes what will the slope of the perpendicular line be if the slope of the line is -0.3?

OpenStudy (amruta1202):

So -0.3 is the slope Now u have the slope of the required line and a point through which it passes

OpenStudy (rayep12):

10/3? i think

OpenStudy (amruta1202):

By slope point form u can easily find the eqn

OpenStudy (amruta1202):

Oops sorry its 10/3

TheSmartOne (thesmartone):

so now we have y = (10/3)x + b we need to plug in the point (3, -8) and find the value of b, the y-intercept

OpenStudy (rayep12):

okay

OpenStudy (amruta1202):

y-y1=m'(x-x1) Where m' is 10/3 the slope of required line and x1 and y1 are the co-ordinates of the given points

OpenStudy (amruta1202):

*point

OpenStudy (rayep12):

so its y=10/3x -18?

TheSmartOne (thesmartone):

correct, or you can use what amruta gave you. That's one step shorter. Not sure which form they want you to give it in.

OpenStudy (amruta1202):

👍🏻